### v1

# UCM-Net: A Lightweight and Efficient Solution for Skin Lesion Segmentation using MLP and CNN

Abstract
Skin cancer is a significant public health problem, and computer-aided diagnosis can help to prevent and treat it. A crucial step for computer-aided diagnosis is accurately segmenting skin lesions in images, which allows for lesion detection, classification, and analysis. However, this task is challenging due to the diverse characteristics of lesions, such as appearance, shape, size, color, texture, and location, as well as image quality issues like noise, artifacts, and occlusions. Deep learning models have recently been applied to skin lesion segmentation, but they have high parameter counts and computational demands, making them unsuitable for mobile health applications. To address this challenge, we propose UCM-Net, a novel, efficient, and lightweight solution that integrates Multi-Layer Perceptions (MLP) and Convolutional Neural Networks (CNN). Unlike conventional UNet architectures, our UCMNet-Block reduces parameter overhead and enhances UCM-Net’s learning capabilities, leading to robust segmentation performance. We validate UCM-Net’s competitiveness through extensive experiments on isic2017 and isic2018 datasets. Remarkably, UCM-Net has less than 50KB parameters and less than 0.05 Giga-Operations Per Second (GLOPs), setting a new possible standard for efficiency in skin lesion segmentation. The source code will be publicly available.
Keywords Medical image segmentation  $\cdot$ Light-weight model  $\cdot$ Mobile health
1 Introduction
Skin cancer poses a significant global health concern and stands as one of the leading cancer types worldwide. Skin cancer can be broadly categorized into two types: melanoma and non-melanoma. While melanoma accounts for only 1% of cases, it is responsible for the majority of deaths due to its aggressive nature. In 2022, it was estimated that melanoma would account for approximately 7,650 deaths in the United States, affecting 5,080 men and 2,570 women [1, 2]. In addition, it is estimated that the United States will have 97,610 new cases of melanoma in 2023. Current statistics suggest that one in five Americans will develop skin cancer at some point in their lives, underscoring the gravity of this issue. Over the past few decades, skin cancer has emerged as a substantial public health problem, resulting in annual expenses of approximately \$ 8.1 billion in the United States alone [3].
Skin cancer [4] is a prevalent and potentially life-threatening disease affecting millions worldwide. Among the various types of skin cancer, malignant melanoma is known for its rapid progression and high mortality rate if not detected and treated early. Early and accurate diagnosis is, therefore, critical to improving patient outcomes. Medical imaging [5], particularly dermatoscopy and dermoscopy, is crucial in diagnosing skin cancer. Dermatologists and healthcare professionals rely on these imaging techniques to examine and analyze skin lesions for signs of malignancy. However, the manual interpretation of such images with a naked eye is a time-consuming and error-prone process, heavily reliant on the expertise of the examining physician [6, 7].
To address these challenges and improve the accuracy and efficiency of skin cancer diagnosis, computer-aided tools and artificial intelligence (AI) have been leveraged in recent years[8, 9, 10]. Skin cancer segmentation, a fundamental step in the diagnostic process, involves delineating the boundaries of skin lesions within medical images. This task is essential for quantifying lesion characteristics, monitoring changes over time, and aiding in the decision-making process for treatment. Segmenting skin lesions from images faces several key challenges [11]: unclear boundaries where the lesion blends into surrounding skin; illumination variations that alter lesion appearance; artifacts like hair and bubbles that obscure lesion boundaries; variability in lesion size and shape; different imaging conditions and resolutions; age-related skin changes affecting texture; complex backgrounds that hinder segmentation; and differences in skin color due to race and climate. Figure 1 presents some representative samples of complex skin lesion.
Figure 1: Complex skin lesion samples
Overcoming these difficulties is crucial for accurate segmentation to enable early diagnosis and treatment of malignant melanoma. Recently, a groundbreaking transformation in skin cancer segmentation has been driven by the development of advanced deep-learning algorithms [12, 13, 14, 15, 16]. These AI-driven approaches have exhibited remarkable capabilities in automating the segmentation of skin lesions, significantly reducing the burden on healthcare professionals and potentially improving diagnostic accuracy. In addition, the rapid advancements in AI techniques and the widespread adoption of smart devices,such as the point-of-care ultrasound (POCUS) devices or smartphones [17, 18, 19], have brought about transformative changes in the healthcare industry [20]. Figure 2 briefly presents the entire diagnose of skin cancer detection with portable devices and AI techniques.
Figure 2: AI diagnose process of skin cancer detection
Patients now have greater access to medical information, remote monitoring, and personalized care, leading to increased satisfaction with their healthcare experiences. However, amidst these advancements, there are still challenges that need to be addressed. One such challenge is the accurate and efficient segmentation of skin lesions for diagnostic purposes within limited computation hardwares and devices. Most of AI medical methods are developed based on deep-learning [21]. The major deep-learning methods utilize expensive computation overhead and a large number of learning parameters to achieve a good prediction result. It is a challenge to embed these methods to hardware-limit devices [22, 23]. In this study, we introduce UCM-Net, a lightweight and robust approach for skin lesion segmentation. UCM-Net leverages a novel hybrid module that combines Convolutional Neural Networks (CNN) and Multi-Layer Perceptrons (MLP) to enhance feature learning while reducing parameters. Utilizing group loss functions, our method surpasses existing machine learning-based techniques in skin lesion segmentation.
Key contributions of UCM-Net include:
  1. 1.
    Hybrid Module: We introduce the UCM-Net Block, a hybrid structure combining CNN and MLP with superior feature-learning capabilities and reduced computation and parameters.
  2. 2.
    Efficient Segmentation: UCM-Net is developed based on UCM-Net Blocks and the base model U-Net, offering a highly efficient method for skin lesion segmentation. It is the first model with less than 50 KB parameters and less than 0.05 Giga-Operations Per Second (GLOPs). UCM-Net is 1177 times faster and has 622 times fewer parameters than U-Net. Compared to the state-of-the-art EGE-UNet, UCM-Net reduces parameter and computation costs by 1.06x and 1.56x.
  3. 3.
    Improved Segmentation: UCM-Net’s segmentation performance is evaluated using mean Intersection over Union (mIoU) and mean Dice similarity score (mDice). On the Isic2017 and Isic2018 datasets, UCM-Net enhances the baseline U-Net model by an average of 3.03% in mIoU and 1.72% in mDice. Notably, UCM-Net outperforms the state-of-the-art EGE-UNet on ISIC 2017 and ISIC 2018 datasets, with respective mean IoU scores of 81.43% (UCM-Net) vs 80.95% and 80.71% (UCM-Net) vs 80.00%.
Figure 3: This figure shows the visualization of comparative experimental results on the ISIC2017 dataset. The X-axis represents mDice score (higher is better), while Y-axis represents mIoU (higher is better). The color depth represents the number of parameters (blue is better).
2 Related works
AI Method Categories and Applications AI-driven approaches for biomedical images can be broadly classified as supervised learning methods, semi-supervised learning methods, and unsupervised learning methods [24, 25, 26]. Supervised learning is a solution with labeled image data, expecting to develop predictive capability. The labeled image data can be the previous patients’ diagnosed results, such as computed tomography(CT), with Clinicians’ analysis. With the labeled data, AI-driven solutions can be developed and performed against the ground truth results. Supervised learning solutions are widely applied to disease classification, tumor location detection, and tumor segmentation [27]. Relatively, unsupervised learning is a discovering process, diving into unlabeled data to capture hidden information. Unsupervised learning solutions derive insights directly from unlabeled medical data without inadequate or biased human supervisions and can be used for information compression, dimensional reduction, super resolution for medical image and sequence data detection and analysis such as protein, DNA and RNA [28]. In recent years, semi-supervised learning is becoming popular, which utilizes a large number of unlabeled data in conjunction with the limited amount of labeled data to train higher-performing models. Semi-supervised learning solutions can be also applied into disease classification and medical segmentation [29, 30].
Supervised Methods of Segmentation As the technique evolves and develops, the solution of AI for medical image segmentation is from purely applying a convolution neural network(CNN) such as U-Net and Att-UNet [31] to a hybrid structure method like TransFuse [32] and SANet [33]. U-Net is a earest CNN solution on biomedical image segmentation, which replaces pooling operators by upsampling operators. Att-UNet is developed on the top of U-Net adding attention structures. TransFuse is a novel approach that combines Transformers and CNNs with late fusion for medical image segmentation, achieving a strong performance while maintaining high efficiency, with potential applications in various medical-related tasks. SANet [34], the Shallow Attention Network, addresses challenges in polyp segmentation by mitigating color inconsistencies, preserving small polyps through shallow attention modules, and balancing pixel distributions, achieving remarkable performance improvements. Swin-UNet is a UNet-like pure Transformer for medical image segmentation that proposed shifted windows as the encoder to extract context features, and a transformer-based decoder with patch expanding layer performs the up-sampling operation to restore the spatial resolution of the feature maps. MedT [35] is also a transformer-based network architecture, a gated axial-attention model that introduces an additional control mechanism in the self-attention module. ConvUNeXt [36], an efficient model inspired by ConvNeXts [37] and based on the classic UNet architecture, achieving excellent medical image segmentation results with a significantly reduced parameter count while incorporating features such as large convolution kernels, depth-wise separable convolution, residual connections, and a lightweight attention mechanism. UNeXt [38] is introduced as an efficient Convolutional multilayer perceptron (MLP) based network that reduces parameters and computational complexity while achieving superior segmentation performance through tokenized MLP blocks and channel shifting, making it suitable for point-of-care applications. MALUNet [39] and its extended version EGE-UNet [40] develop new attention modules to significantly reduce parameters and computational complexity while achieving powerful skin lesion segmentation performance, making it highly suitable for resource-constrained clinical environments.
Figure 4: UCM-Net Structure
3 UCM-Net
Network Design Figure 4 provides a comprehensive view of the structural framework of UCM-Net, an advanced architecture that showcases a distinctive U-Shape design. Our design is developed from U-Net. UCM-Net includes a down-sampling encoder and an up-sampling decoder, resulting in a high-powered network for skin lesion segmentation. The entirety of the network encompasses six stages of encoder-decoder units, each equipped with channel capacities of {8, 16, 24, 32, 48, 64}. Within each stage, we leverage a convolutional block alongside our novel UCMNet block, facilitating the extraction and acquisition of essential features. In the convolutional block, we opt for a kernel size of 1, a choice that serves to further curtail the parameter count. Our innovative UCMNet introduces a hybrid structure module, wherein an amalgamation of a Multi-Layer Perceptron (MLP) linear component and Convolutional Neural Network (CNN) is employed, bolstered by the inclusion of skip connections. This strategic amalgamation fortifies the network’s prowess in feature acquisition and learning capabilities.
Convolution Block In our designing, we contain two different convolution blocks. The difference between conv block 1 and conv block 2 is from the number of kernel size. Small number of kernel size can not only reduce the number of learning weights, but also reduce the computations usage [41]. We set conv block 2’s kernel to 1 $\times$ 1. And we maintain conv block 1’s kernel size to 3 $\times$ 3 normally since fewer the beginning convolution kernel can affect the features entropy performance from input.
# Input: X,the feature map with shape [Batch(B), Channel(C), Height(H), Width(W)]
# Output: Out,the feature map with shape [B, Height*Width(N),C]
# Operator: Conv, 2D Convolution  LN, LayerNorm  BN, BatchNorm,  Linear, Linear Transformation  Leaky, Leaky RelU
# UCM-Net Block Processing Pipeline
B, C, H, W = X.shape()
# Transform Feature from [B,C,H,W] to [B,H*W,C]
X = X.flatten(2).trnaspose(1,2)
# Copy feature for later residual addition
X1 = copy(X)
X = Linear(LN(X))
B, N, C = X.shape()
# Transform Feature from [B,H*W,C] to [B,C,H,W]
X = X.transpose(1,2).view(B,C,H,W)
X = Conv(LN(X))
X = Conv(Leaky(X))
# Transform Feature from [B,C,H,W] to [B,H*W,C]
X = X.flatten(2).trnaspose(1,2)
X = Linear(BN(X))
# Transform Feature from [B,H*W,C] to [B,C,H,W]
X = X.transpose(1,2).view(B,C,H,W)
X = Conv(LN(X))
# Transform Feature from [B,C,H,W] to [B,H*W,C]
X = X.flatten(2).trnaspose(1,2)
# Output with residual addition
Out = X + X1
Algorithm 1 PyTorch-style pseudocode for UCM-Net Block
UCM-Net Block The pseudocode of UCM-Net Block 1 presents our defined sequence of operations, which is how we combine CNN with MLP(Linear transformation operation) for feature learning. The input feature structure for CNN operation is a four-dimensional structure containing batch, channel, height, and weight. However, the input feature structure for MLP is a three-dimensional structure, which includes batch, channel, and vector. As the pseudocode shows, we implement several feature transformations to service CNN and MLP learning in different layers.
Loss functions In our solution, we selected the group loss function from EGE-UNet [40]. The loss function can calculate the loss from the scaled layer masks in different stages with ground truth masks. Equation 1 and 2 present the stage loss in different stage layer and output loss in the output layer, which calculated by binary cross-entropy (BCE) and dice loss (Dice) components.
$Loss_{Stage}=\text{Bce}(StagePred,Target)+\text{Dice}(StagePred,Target)$ (1)
$Loss_{Output}=\text{Bce}(OutputPred,Target)+\text{Dice}(OutputPred,Target)$ (2)
Equation 3 represents the loss in different stages. Equation 4 presents the group loss that includes different stages loss and output loss. $\lambda_{i}$ is the weight for different stage. In this paper, we set $\lambda_{i}$ to 0.1, 0.2, 0.3, 0.4 and 0.5 based on i-th stage as shown in Figure 1.
$Loss_{\text{Stages}}=\sum_{i=1}^{5}\lambda_{i}\times Loss_{\text{Stagei}}$ (3)
$GroupLoss$ $=Loss_{\text{Output}}+Loss_{\text{Stages}}$ (4)
4 Experiments and Results
4.1 Experiments Setting
Datasets To evaluate the efficiency and performance of model with other published models, we pick the two public skin segementation datasets from International Skin Imaging Collaboration, namely ISIC 2017 and ISIC2018. we select two public skin lesion segmentation datasets, namely ISIC2017 [42, 43] and ISIC2018 [44, 45]. The ISIC2017 dataset comprises 2150 dermoscopy images, and ISIC2018 includes 2694 images. We noted that earlier studies[40, 39] have already presented a dataset version with a pre-established train-test partition, maintaining a 7:3 ratio. In our experimental setup, we opted to utilize the previously published dataset version.
Implementation Details Our UCM-Net is implemented with Pytorch [46] framework. All experiments are conducted on the instance node at Lambda [47] that has a single NVIDIA RTX A6000 GPU (24 GB), 14vCPUs, 46 GiB RAM and 512 GiB SSD. The images are normalized and resized to 256$\times$256. Simple data augmentations are applied, including horizontal flipping, vertical flipping, and random rotation. We noticed the prior studies [40, 39] applied initial image processing with the calculated mean and standard deviation (std) values of the whole train and test datasets separately. While this approach can potentially enhance their models’ training and testing performance, the outcomes are notably influenced by the computed mean and std values. Additionally, if the test dataset’s context information is unknown, this operation can render the trained model less practical. In our experiment, we don’t calculate the mean and std values based on the train and test datasets. Besides, TransFuse-S and SwinNet require the pre-train models with the specified input image size in their encoding stage. To enable fair benchmark testing, we follow the image input size for the TransFuse-S [32, 48] and SwinNet [33, 49] to 192$\times$256 and 224$\times$224, Correspondingly. For the optimizer, we select AdamW [50] initialized with a learning rate of 0.001 and a weight decay of 0.01. The CosineAnnealingLR [51] is Utilized as the scheduler with a maximum number of iterations of 50 and a minimum learning rate of 1e-5. A total of 300 epochs are trained with a training batch size of 8 and a testing batch size of 1.
Evaluate Metrics To assess the predictive performance of our methods, we employ mean Intersection over Union (mIoU) and mean Dice similarity score (mDice) as evaluation metrics. It’s worth noting that previous studies [40, 39] and [38] have employed distinct calculation methods for mIoU and mDice. To comprehensively compare the performance predictions, our experiments include the presentation of mIoU, mDice, mIoU*, and mDice* results. These results are calculated using the following equations:
$\text{IoU}=\frac{\text{intersection}}{\text{union}}$ (5)
Dice=2×intersectionsum of pixels in prediction+sum of pixels in ground $\text{Dice}=\frac{2\times\text{intersection}}{\text{sum of pixels in prediction}+\text{sum of pixels in ground truth}}$ (6)
where intersection represents the number of pixels that are common between the predicted output and the ground truth, and union represents the total number of pixels in both the predicted output and the ground truth.
$\text{mIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{i}$ (7)
$\text{mDice}=\frac{1}{N}\sum_{i=1}^{N}\text{Dice}_{i}$ (8)
where N is the number of images, $\text{IoU}_{i}$ represents the IoU score for image i and $\text{Dice}_{i}$ represents the Dice score for image i.
[…]
where TP represents the number of true positive pixels, FP represents the number of false positive pixels and FN represents the number of false negative pixels.
[…]
where $\text{TP}_{sum}$ represents the total number of true positive pixels for images, $\text{FP}_{sum}$ represents the total number of false positive pixels for images and $\text{FN}_{sum}$ represents the total number of false negative pixels for images.
In our benchmark experiments, we evaluate our method’s performance and compare the results among other published efficient models’. To ensure a fair comparison, we perform three sets of experiments for each method and subsequently present the mean and std of the prediction outcomes across each dataset.
4.2 Performance Comparisons
Table 1: Comparative prediction results on the ISIC2017 and ISIC2018 dataset
Dataset Models Year mIoU(%)↑ mDice(%)↑ mIoU*(%)↑ mDice*(%)↑
isic2017 U-Net [52, 53] 2015 78.12 ± 0.175 85.97 ± 0.196 76.42 ± 0.381 86.63 ± 0.245
Att-UNet [54, 31] 2018 78.45 ± 0.113 86.22 ± 0.124 77.14 ± 0.097 87.10 ± 0.062
${}^{\ast}$ [32, 48] 2021 79.45 ± 0.427 87.12 ± 0.402 78.98 ± 1.222 88.25 ± 0.762
${}^{\ast}$[34, 55] 2021 79.34 ± 0.139 86.94 ± 0.155 78.64 ± 0.295 88.04 ± 0.185
${}^{\ast}$ [33, 49] 2021 80.51 ± 0.152 87.81 ± 0.157 80.48 ± 0.511 89.19 ± 0.313
MedT [35, 56] 2021 79.05 ± 0.231 86.59 ± 0.125 77.61 ± 0.121 87.40 ± 0.401
ConvUNeXt [36, 57] 2022 78.78 ± 0.362 86.21 ± 0.267 76.98 ± 0.490 86.99 ± 0.313
UNeXt-S [38, 58] 2022 78.79 ± 0.234 86.30 ± 0.140 77.48 ± 0.466 87.31 ± 0.296
MALUNet [39, 59] 2022 79.97 ± 0.389 87.23 ± 0.345 79.11 ± 0.345 88.34 ± 0.215
EGE-UNet [40, 60] 2023 80.00 ± 0.010 87.07 ± 0.053 79.26 ± 0.028 88.43 ± 0.017
UCM-Net (ours) 2023 80.71 ± 0.345 87.66 ± 0.221 79.29 ± 0.188 88.45 ± 0.117
isic2018 U-Net [52, 53] 2015 79.86 ± 0.075 87.57 ± 0.085 78.27 ± 0.300 87.81 ± 0.188
Att-UNet [54, 31] 2018 80.05 ± 0.079 87.62 ± 0.078 78.38 ± 0.151 87.88 ± 0.095
${}^{\ast}$ [32, 48] 2021 81.20 ± 0.049 88.42 ± 0.025 80.61 ± 0.463 89.26 ± 0.284
${}^{\ast}$ [34, 55] 2021 80.37 ± 0.124 87.87 ± 0.114 79.39 ± 0.135 88.51 ± 0.084
${}^{\ast}$ [33, 49] 2021 81.41 ± 0.069 88.58 ± 0.019 80.72 ± 0.069 89.33 ± 0.042
MedT [35, 56] 2021 80.34 ± 0.034 87.77 ± 0.107 79.29 ± 0.411 88.45 ± 0.251
ConvUNeXt [36, 57] 2022 80.51 ± 0.043 87.99 ± 0.049 78.71 ± 0.128 88.09 ± 0.080
UNeXt-S [38, 58] 2022 80.70 ± 0.226 88.17 ± 0.194 79.26 ± 0.497 88.43 ± 0.309
MALUNet [39, 59] 2022 80.95 ± 0.393 88.25 ± 0.315 79.99 ± 0.644 $8$ 0.398
EGE-UNet [40, 60] 2023 80.95 ± 0.054 88.23 ± 0.096 80.11 ± 0.221 88.95 ± 0.137
UCM-Net (ours) 2023 81.43 ± 0.020 88.53 ± 0.024 80.74 ± 0.285 89.34 ± 0.175
  • ${}^{\ast}$: this method needs the pre-train model on training.
Table 2: Comparative performance results on models’ computations and the number of parameters.
Models Year Image Size (H x W) Params↓ GFLOPs↓
U-Net [52, 53] 2015 256 x 256 31,037,633 54.7378
Att-UNet [54, 31] 2018 256 x 256 34,878,573 66.6318
TransFuse-S [32, 48] 2021 192 x 256 26,248,725 8.6462
SANet [34, 55] 2021 256 x 256 23,899,497 5.9983
SwinNet [33, 49] 2021 224 x 224 20,076,204 5.5635
MedT [35, 56] 2021 256 x 256 1,564,202 2.4061
ConvUNeXt [36, 57] 2022 256 x 256 3,505,697 7.2537
UNeXt-S [38, 58] 2022 256 x 256 253,561 0.1038
MALUNet [39, 59] 2022 256 x 256 177,943 0.0830
EGE-UNet [40, 60] 2023 256 x 256 53,374 0.0721
UCM-Net(ours) 2023 256 x 256 49,932 0.0465
Figure 5: Vision performance comparison on samples
Table 1 comprehensively evaluates the performance of our UCM-Net, a novel skin lesion segmentation model, compared to well-established models, using the widely recognized ISIC2017 and ISIC2018 datasets. Introduced in 2023, UCM-Net is a robust and highly competitive solution in this domain. One of the key takeaways from the table is UCM-Net’s ability to outperform EGE-UNet, which had previously held the title of the state-of-the-art model for skin lesion segmentation. Our model achieves superior results across various prediction metrics, emphasizing its advancement in the field and its potential to redefine the standard for accurate skin lesion delineation. Moreover, UCM-Net’s performance is notably competitive even when compared to SwinNet, a model that relies on pre-trained models during training. Table 2 complements this assessment by comparing computational aspects and the number of parameters for various segmentation models. Remarkably, UCM-Net, operating with the same number of channels {8,16,24,32,48,64} and image size, as EGE-UNet, boasts fewer parameters and lower GFLOPs. Additionally, even when compared to TransFuse-S and SwinNet, which operate with smaller image sizes, UCM-Net demonstrates faster computational speed. In Figure 5, we present a visual exhibition of all the models’ segmentation outputs. This figure directly compares our segmentation results, those produced by other methods, and the ground truth, all displayed side by side using representative sample images. Notably, our segmentation results demonstrate a remarkable level of accuracy, closely resembling the ground truth annotations. Tables 1-2 and Figure 5 collectively underscore UCM-Net’s exceptional performance and efficiency in skin lesion segmentation, affirming its potential to make a substantial impact in advancing early skin cancer diagnosis and treatment.
4.3 Ablation results
To demonstrate the efficiency and effectiveness of our proposed modules, we conducted a series of ablation experiments on dataset ISIC2017. We develop UCM-Net based on U-Net. Figure 6 shows the different block structures in the stages among the compared models. Table 3 shows the ablation experiments’ results including the number of parameters, Giga Flops, mIoU score and mDice score. U-Net variant and variant 1 are six-stage U-Nets with the stage channels {8,16,24,32,48,64}. The details of UCM-Net are illustrated in Figure 4.
Figure 6: Stage Block Structures in ablation experiments.
A: U-Net and U-Net Variant. B: U-Net Variant 1. C: UCM-Net
Figure 6: Stage Block Structures in ablation experiments.
Table 3: Ablation experiments’ results on the ISIC2017 dataset
Models Structure Reference Params↓ GFLOPs↓ mIoU(%)↑ mDice(%)↑
U-Net(baseline) Figure 6 (A) 310,376,33 54.7378 78.34 86.23
U-Net Variant Figure 6 (A) 248,531 0.5715 78.48 86.22
U-Net Variant 1 Figure 6 (B) 148,157 0.3700 73.89 82.36
UCM-Net Figure 6 (C) 49,932 0.0465 79.76 86.94
UCM-Net + Group Loss Figure 6 (C) 49,932 0.0465 80.63 87.64
Figure 7: IoU vs Epoch results of ablation experiments
Figure 8: Dice vs Epoch results of ablation experiments
As the results of the U-Net and U-Net variants are shown in Table 3, although the number of parameters is reduced, the U-Net variant performs better with the six-stage structure. When we set the one convolution kernel to 1 to reduce the number of parameters, the model’s performance drops severely. However, when we replaced the convolution with our proposed UCMNet block, the results showed that the model’s performance improved significantly. The UCM-Net, as depicted in Figure 4 and Figure 6(C), with 49,932 parameters and 0.0465 GFLOPs, outperforms the U-Net variant with 248,531 parameters and the baseline U-Net with 31,037,633 parameters in terms of both mean Intersection over Union (mIoU) and mean Dice Similarity Coefficient (mDice) metrics.
Furthermore, when incorporating the Group Loss into the UCM-Net architecture, denoted as "UCM-Net + Group Loss", the model’s performance continued to excel. This enhancement resulted in a higher mIoU of 80.63% and a mDice of 87.64%, demonstrating the effectiveness of the proposed Group Loss in further improving segmentation accuracy. Figures 7-8 show that "UCM-Net + Group Loss" always presents the high scores of mIoU and mDice with the training epoch increase.
The above findings in the ablation experiments underscore the significance of architectural innovations such as the UCMNet block in achieving superior semantic segmentation performance, even with fewer parameters and computational complexity than the U-Net baseline.
5 Conclusion
This paper introduces UCM-Net, a novel, lightweight, and highly efficient solution. UCM-Net combines MLP and CNN, providing robust feature learning capabilities while maintaining a minimal parameter count and reduced computational demand. We applied this innovative approach to the challenging task of skin lesion segmentation, conducting comprehensive experiments with a range of evaluation metrics to showcase its effectiveness and efficiency. The results of our extensive experiments unequivocally demonstrate UCM-Net’s superior performance compared to the state-of-the-art EGE-UNet. Remarkably, UCM-Net is the first model with fewer than 50KB parameters and consuming less than 0.05 GLOPs for skin lesion segmentation. Looking forward to future research endeavors, we aim to expand the application of UCM-Net to other critical medical image tasks, advance the field, and explore how this efficient architecture can contribute to a broader spectrum of healthcare applications, potentially revolutionizing how we utilize deep learning for medical image analysis.